OpenStudy (samsan9):

what is the radius of a circle with the equation x^2+y^2+2x+4y-9=0?

OpenStudy (zzr0ck3r):

\(x^2+2x+y^2+4y-9=0\\(x+1)^2-1+(y+2)^2-4-9=0\\(x+1)^2+(y+2)^2=14\)

OpenStudy (samsan9):

so the radius is 14?

OpenStudy (zzr0ck3r):

now use the equation of a circle to get the radius I completed the square two times...

OpenStudy (anonymous):

http://www.purplemath.com/modules/sqrcircle.htm

OpenStudy (zzr0ck3r):

what is the equation of a circle?

OpenStudy (anonymous):

check that link

OpenStudy (samsan9):

(x+h)^2+(y-k)^2=r^2

OpenStudy (zzr0ck3r):

so we have r^2 = 14 what is r?

OpenStudy (anonymous):

\[\sqrt{14}\]

OpenStudy (zzr0ck3r):

it is important that you see how I did this problem, do you?

OpenStudy (samsan9):

i guess its sqrt(14)

OpenStudy (samsan9):

or 196

OpenStudy (zzr0ck3r):

what is 196?

OpenStudy (samsan9):

nevermind i guess its square root of 14 but it also says to put the answer into the nearest thousand form

OpenStudy (samsan9):

so is it 3.472 ?

OpenStudy (zzr0ck3r):

if thats what your calculator says...